If I could get some help with the following question that would be great. Please include explanations if possible.

Suppose that a firm is currently employing 20 workers, the only variable input, at a wage rate of $60. The average product of labor is 30, the last worker added 12 units to total output, and total fixed cost is $3600.  

a. what is marginal cost?

b. what is average variable cost?

c. how much output is being produced?

d. what is average total cost?

e. is average variable cost increasing, constant, or decreasing? what about average total cost?
Answer #1

a) MC = change in TC/change in output

60/12 = 5

b) AVC = TVC/output

= (20*60)/(30*20)

= 1200/600

= 2

c) Output being produced = APL*Labor = 30*20 = 600

d) ATC = TC/output

TC=TFC+TVC

= (3600+1200)/600

= 8